# Tidal disruption of a neutron star near naked singularity

### v2

@@ Abstract
We investigate the tidal disruption of a neutron star (NS) near a black hole (BH), and for the first time, to the best of our knowledge, near a naked singularity (NaS). For a BH with a mass greater than about $10M_{\odot}$, the tidal disruption of NS should occur within the event horizon, and hence neither can the stellar material escape nor a distant observer observe the disruption. Since NaS does not have an event horizon, a significant portion of the NS’s material can escape, and the tidal disruption can be observed by a distant observer. One could identify such an event from the observed emission from the disrupted NS’s material and the decay of the light curve of the disruption event. The escape of a significant fraction of the NS’s material may also have implications for the heavy elements in the universe. Moreover, observing such an event can be useful for confirming a NaS, probing its spacetime, and studying the motion of matter in such a geometry. This may help constrain the NS parameters and equation of state models. As a first step in this direction, we calculate here the tidal disruption radius and other parameters for a specific type (Joshi-Malafarina-Narayan type 1) of NaS and compare our results with observations.
@@ I Introduction
In recent decades, tidal disruption events have been the subject of much research. Rees’s early theoretical work Rees:1988 predicted that supermassive black holes would disrupt stars, with a portion of the stellar debris falling back to fuel luminous flares. Several TDE candidates in the X-ray and UV bands were later confirmed by observational campaigns Pahari:2024rev ; Cocchi:2023mgm ; well-researched examples, like ASASSN-14li, provide comprehensive multi-wavelength data Holoien:2016 . The sample of known TDEs has been greatly increased in recent years by large optical transient surveys (e.g., ZTF, ASAS-SN), which have shown a greater diversity in luminosity, emission features, and timescales Komossa:2015 ; Dai:2018 . Finding faint and rapidly evolving TDEs, like ASASSN-23bd, further raises the possibility of new subclasses and demonstrates how TDEs can be used to study accretion physics and black hole demographics Hoogendam:2024tnk ; Bhattacharya:2025qps ; Wevers:2023ksb ; Zhou:2025lzg ; Pasham:2022oee ; Bhattacharyya:2017tos ; Ozel:2016oaf .
@@
According to Lattimer:1976kbf , BHs with masses between 8-17 solar masses have enough mass to trigger the typical neutron star (NS) disruption that an asymptotic observer could see; the exact range depends on the NS equation of state. A binary NS, a BH-NS, and a NaS-NS binary system coalesce, and this phenomenon is important because it results in high-energy events, such as gamma-ray bursts (GRBs). We now know that short-duration (sGRB) events, a subtype of GRB signals, originate from the cataclysmic merger of binary NSs. The gravitational-wave signal GW170817, which was generated when two NSs spiraled together and collided, was observed by scientists for the first time on August 17, 2017 LIGOScientific:2017vwq . This event was unique among BH merger findings, as it was also observed in light, initially as a GRB and later throughout the entire electromagnetic spectrum. For the first time, scientists were able to observe the same cosmic event (kilonova) using both conventional telescopes and gravitational waves. This discovery demonstrated how heavy elements like gold and platinum are created and showed that short gamma-ray bursts are powered by merging NSs Metzger:2019zeh ; Abbott:2017multimessenger ; Abbott:2017kilonova ; Abbott:2019properties ; Abbott:2020gwtc2 .
@@
It should be noted that the likelihood of observing tidal disruption of NSs by supermassive NaSs is dependent on the pulsar or NS population close to galactic centers, in addition to the presence or population of supermassive NaSs. For instance, our Galactic Center has a ”missing pulsar problem” because hardly any pulsars have been discovered there. There may be fewer pulsars close to a galactic center because of the high density of dark matter (DM), which may cause DM particles to gather at the core of the NS and create a tiny BH or NaS, which then accretes and transforms the entire NS into a BH or NaS. Therefore, the absence of NaS may not be indicated by the non-observation or low likelihood of NS disruption close to galactic centers. Therefore, the absence of NSs near galactic centers or the low likelihood of their disruption may indicate the absence of NSs rather than the absence of NaS. Adarsha . In Schwarzschild spacetime, the innermost stable circular orbit (ISCO) is located at $r=6GM/c^{2}$. According to the Novikov-Thorne thin-disk accretion model, a particle in this orbit has a specific energy of $e=\sqrt{8/9}$, which translates to a radiative efficiency of roughly $5.7\%$. In contrast, the conserved energy per unit rest mass becomes $e=1$ at $r=4GM/c^{2}$, which is the marginally bound (and unstable) circular orbit. Theoretically, particles around $r=2GM/c^{2}$ could, in an idealized sense, reach efficiencies approaching $100\%$ if accretion were extremely non-circular and matter could lose almost all of its energy, for example, during extreme tidal disruption events or dissipative infall. However, since radiation generated close to the horizon cannot escape to infinity, such a complete conversion of rest-mass energy is not possible in realistic Schwarzschild accretion. On the other hand, in the case of the NaS, the accretion efficiency is always $100\%$. Here we consider the Joshi-Malafarina-Narayan (JMN) 1 NaS psJoshi1 .
The line element of a general spherically symmetric, static spacetime is given as follows (we discuss it in detail in Appendix A):
@@
We investigate the effect of tidal forces in a static spherically symmetric (SSS) spacetime. To explore tidal forces in the framework of general relativity, we analyze the geodesic deviation equation:
@@
where $R^{\mu}_{\nu\rho\sigma}$ and $v^{\nu}$ denote the Riemann curvature tensor and the unit tangent vector of the geodesic, respectively, and $\eta^{\mu}$ represents the geodesic deviation vector. In the presence of strong energy conditions and non-zero curvature, each point on the test body follows a unique geodesic, leading to stretching and/or squeezing—known as tidal effects. To quantify these effects, which measures the distance between infinitesimally close geodesics. The tetrad basis related to the freely falling frame is expressed as:
@@
where $\eta_{\hat{\mu}\hat{\nu}}=\text{diag}(-1,1,1,1)$ represents components of the Minkowski metric. The hat indices denote the tetrad basis, while those without denote the coordinate basis. Additionally, $\hat{e}^{\mu}_{0}=v^{\mu}$, implying that the tetrad basis equals the 4-velocity vector of the observer when timelike Lima Junior et al. (2022). The unit vectors $\{\hat{e}^{\mu}_{1},\hat{e}^{\mu}_{2},\hat{e}^{\mu}_{3}\}$ correspond to orthogonal spatial directions in the observer’s frame chandrasekhar1983mathematical . Using the tetrad basis from Eq. (7), the geodesic deviation vector (or separation vector) can be expanded as:
@@
The different Keplerian orbits tend to be compressed and stretched as the star approaches the BH, causing instability in the hydrostatic equilibrium. The specific energy distribution is then widened by pressure forces that transfer energy within the star. A significantly broader dispersion of internal energies characterizes the star following the encounter, with some fluid having a positive energy (and thus staying unbound) and some having a negative energy (and hence being bound to the BH). The fluid elements move in Keplerian orbits. Following a Keplerian period T, the bound debris return near the pericenter and are connected to their (negative) energy E by:
@@
The second essential premise that we must make in order to produce the “standard” $t^{-5/3}$ light curve, that is, the energy distribution ($dm/dE$), is uniform, independent of time $T$. Rees (1988) merely assumed this to be the simple case. Evans and Kochanek’s (1989) Evans:1989qe , numerical simulations later seemed to display a uniform energy distribution, indicating that the light curve is roughly proportional to $t^{-5/3}$. In Lodato:2008fr discuss that the energy distribution need not be uniform and instead depends on the star’s characteristics, particularly its interior structure, like the density distribution within the star. One distinctive feature of this theory is the time dependency of the light curve of such tidal events, which is represented as L(t) proportional to $t^{-5/3}$. Since then, a $t^{-5/3}$ light curve has been frequently used to fit the observed luminosities of occurrences thought to be star disturbances. Hence, some observational study Jankovic:2023nyy ; Coughlin:2019pqk
@@ II Tidal disruption radius in BHs and naked singularities
The way stars are destroyed close to supermassive compact objects—which might not be typical black holes but instead have scalar ”hair” or even be naked singularities—is examined in Andre:2024bia . The authors demonstrate how scalar fields can significantly increase tidal forces in the vicinity of black holes. At the same time, flares may be powered by stellar collisions with a dense ”gray shell” in the case of a naked singularity. According to their findings, tidal disruption events might provide a means of differentiating between black holes and other, more unusual central objects in galaxies. The tidal disruption radius is calculated on the principle of the radial tidal force exerted on the test particle, Eq. (12)(here we consider a NS with some solar mass), and the gravity on the surface of the test particle is equal. However, assuming that the tidal disruption of the star takes place when the self-gravitational attraction on its surface equals the tidal force, which is an elementary but useful approximation for the tidal disruption radius,
$\mathinner{\!\left\lvert T_{R}\right\rvert}R_{n}=\frac{Gm}{R_{n}^{2}}.$ (17)
Here, the radial tidal force component is,
@@ II.2 Roche limit in JMN1 NaS spacetime
Fig. (1) demonstrates tidal disruption radius in JMN1 NaS for $M_{0}<2/3$. As shown in Eq. (24), a real positive value of $T_{R}$ is when $M_{0}<2/3$, which implies that tidal disruption is because of the stretching forces while $M_{0}>2/3$ compressive forces are present. Tidal deformation is also closely related to the Jacobi fields that are fully explained in Appendix B.
@@
The evolution of flares with time near a supermassive compact object or the merger of two compact binaries arises due to the disruption of the compact object. Here, we analytically show the difference between naked singularities and BHs, relating the light curve to the internal density distribution of the star. As it has been previously shown Lodato:2008fr , the standard light curve is proportional to $t^{-5/3}$ only holds at late times. Unbound debris always carries larger mechanical energies than the rest mass energy.
@@
As shown in the strong gravity correction order $1/c^{2}$, $dE/d\tau>0$, and also depends on $\tau^{-7/3}$. From Fig.(1), we can conclude that tidal disruption within the ISCO to the event horizon is possible within the stellar mass mergers of binary BH and NS ($1.5M_{\odot}$ to $7.3M_{\odot}$). In the Schwarzschild case, the innermost unstable circular orbit (IUCO) is $4M$. Hence, in between the $4M$ to $6M$ unstable circular orbit can be possible. If the tidal disruption radius is within this range, the tidal disruption radius is the periapsis of the NS. If a NS approaches a BH so closely that it is affected by tidal distortions and even disruption, it cannot be considered a “point mass” interacting with the BH. Such effects become significant when the ‘tidal radius’ is as small as the pericenter, $R_{T}<r_{min}$. Therefore, complete disruption requires $\beta=R_{T}/r_{min}<1$. For JMN-1 NaS proper time for a circular orbit is given as follows:
$\tau=\frac{2\pi r\sqrt{R_{b}}}{\sqrt{GM}}\sqrt{1-\frac{3GM}{c^{2}R_{b}}}.$ (29)

### v3

@@ Abstract
We investigate the tidal disruption of a neutron star (NS) near a black hole (BH) and a naked singularity (NaS). For a BH with a mass greater than about $10M_{\odot}$, the tidal disruption of an NS should occur within the event horizon, and hence neither can the stellar material escape nor a distant observer observe the disruption. Since a NaS does not have an event horizon, a significant portion of the NS’s material can escape, and the tidal disruption can be observed by a distant observer. One could identify such an event from the observed emission from the disrupted NS’s material and the decay of the light curve of the disruption event. The escape of a significant fraction of the NS’s material may also have implications for the heavy elements in the universe. Moreover, the observation of such an event can be useful to confirm a NaS, to probe its spacetime and the motion of matter in such a spacetime, and also to constrain NS parameters and equation of state models. As a first step in this direction, we calculate the tidal disruption radius and other parameters for a specific type (Joshi-Malafarina-Narayan type 1) of NaS and compare our results with observations.
@@ I Introduction
The phenomenon of stellar tidal disruption by massive black holes was first theoretically proposed by Hills (1975) Hills , with observational implications later explored by Lacy et al. (1982) Lacy . Rees (1988) summarized these developments and predicted a fallback rate scaling as $\sim t^{-5/2}$ Rees:1988 . This behaviour was subsequently refined by Phinney (1989) Phinney , who showed that the post-disruption accretion rate follows the now-standard $\sim t^{-5/3}$ scaling rate. Early analytic models of stellar disruption and the formation of bound and unbound debris, in a relativistic framework by Evans and Kochanek Evans:1989qe . Observational progress in identifying and characterizing tidal disruption events (TDEs) has accelerated significantly in recent years, as summarized in the comprehensive review by Gezari (2021) Gezari . The relativistic dependence of the debris energy distribution and the resulting mass fallback rate has been analyzed in detail in Jankovic:2023nyy . In partial disruption scenarios, the mass fallback rate can follow alternative power-law scalings, as demonstrated in relativistic analyses of partial stellar disruption Coughlin:2019pqk . Large samples of optically, UV-, and X-ray-selected TDEs have been compiled in recent catalogues, including those by Hammerstein et al. (2023) Hammerstein , Yao et al. (2023) Yao , and van Velzen et al. (2021) vanVelzen . These events have been discovered primarily through wide-field transient surveys such as the Zwicky Transient Facility (ZTF), the All-Sky Automated Survey for Supernovae (ASAS-SN), the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS), and the Sloan Digital Sky Survey (SDSS), among others. Several TDE candidates in the X-ray and UV bands were later confirmed by observational campaigns Pahari:2024rev ; Cocchi:2023mgm ; well-researched examples, like ASASSN-14li, provide comprehensive multi-wavelength data Holoien:2016 . The sample of known TDEs has been greatly increased in recent years by large optical transient surveys (e.g., ZTF, ASAS-SN), which have shown a greater diversity in luminosity, emission features, and timescales Komossa:2015 ; Dai:2018 . Finding faint and rapidly evolving TDEs, like ASASSN-23bd, further raises the possibility of new subclasses and demonstrates how TDEs can be used to study accretion physics and black hole demographics Hoogendam:2024tnk ; Bhattacharya:2025qps ; Wevers:2023ksb ; Zhou:2025lzg ; Pasham:2022oee ; Bhattacharyya:2017tos ; Ozel:2016oaf .
@@
While black holes represent the most widely accepted outcome of gravitational collapse, general relativity also admits solutions in which spacetime singularities are not hidden behind event horizons. Such naked singularities arise in a variety of collapse scenarios and continue to play a central role in discussions of the cosmic censorship conjecture. Over the years, considerable effort has been devoted to understanding whether these objects can be distinguished observationally from black holes, through phenomena such as gravitational lensing, accretion dynamics, and strong-field signatures Joshi:2008 ; Virbhadra:2002 ; Bambi:2013 . These studies motivate the search for invariant, physically measurable quantities capable of probing the extreme curvature regions associated with horizonless compact objects. According to Lattimer:1976kbf , BHs with masses between 8-17 solar masses have enough mass to trigger the typical neutron star (NS) disruption that an asymptotic observer could see; the exact range depends on the NS equation of state. A binary NS, a BH-NS, and a NaS-NS binary system coalesce, and this phenomenon is important because it results in high-energy events, such as gamma-ray bursts (GRBs). We now know that short-duration (sGRB) events, a subtype of GRB signals, originate from the cataclysmic merger of binary NSs. The gravitational-wave signal GW170817, which was generated when two NSs spiraled together and collided, was observed by scientists for the first time on August 17, 2017 LIGOScientific:2017vwq . This event was unique among BH merger findings, as it was also observed in light, initially as a GRB and later throughout the entire electromagnetic spectrum. For the first time, scientists were able to observe the same cosmic event (kilonova) using both conventional telescopes and gravitational waves. This discovery demonstrated how heavy elements like gold and platinum are created and showed that short gamma-ray bursts are powered by merging NSs Metzger:2019zeh ; Abbott:2017multimessenger ; Abbott:2017kilonova ; Abbott:2019properties ; Abbott:2020gwtc2 .
@@
Relativistic tidal disruption in Schwarzschild spacetime has been extensively studied since the pioneering analyses of Carter and Luminet Carter:1982 ; Carter:1983 ; Luminet:1986 , which established the role of strong-field tidal forces in stellar deformation and debris energetics. Subsequent work has refined these results within fully relativistic frameworks. A key requirement for any robust probe of strong-field gravity is invariance under coordinate transformations. Curvature invariants and covariantly defined tidal indicators satisfy this criterion, encoding spacetime geometry directly through the Riemann tensor and its contractions, and have proven effective in distinguishing black hole geometries even in extreme relativistic regimes Cherubini:2002 ; Abdelqader:2015 . This motivates the use of tidal observables as invariant diagnostics of strong-curvature regions in more general compact-object spacetimes.
It should be noted that the likelihood of observing tidal disruption of NSs by supermassive NaSs is dependent on the pulsar or NS population close to galactic centers, in addition to the presence or population of supermassive NaSs. For instance, our Galactic Center has a “missing pulsar problem” because hardly any pulsars have been discovered there. There may be fewer pulsars close to a galactic center because of the high density of dark matter (DM), which may cause DM particles to gather at the core of the NS and create a tiny BH or NaS, which then accretes and transforms the entire NS into a BH or NaS. Therefore, the absence of NaS may not be indicated by the non-observation or low likelihood of NS disruption close to galactic centers. Therefore, the absence of NSs near galactic centers or the low likelihood of their disruption may indicate the absence of NSs rather than the absence of NaS. Adarsha . In Schwarzschild spacetime, the innermost stable circular orbit (ISCO) is located at $r=6GM/c^{2}$. For a test particle in the Schwarzschild geometry, the specific energy at the innermost stable circular orbit (ISCO), located at $r=6GM/c^{2}$, is $E=2\sqrt{2}/3=\sqrt{8/9}$. Thus, a particle falling from rest at infinity must radiate an energy $(1-2\sqrt{2}/3)mc^{2}\approx 0.057\,mc^{2}$ to reach this orbit, implying an accretion efficiency of $\sim 5.7\%$. These values are determined solely by the orbital dynamics of circular geodesics in the Schwarzschild spacetime and are not specific to any model, which instead provides the detailed radiative structure of the disc.
One might naively expect that material approaching $r=2GM/c^{2}$ could, under ideal circumstances, radiate nearly all of its binding energy, for example if the motion were strongly non–circular and significant dissipation occurred during tidal interactions or infall. In reality, for a Schwarzschild black hole, radiation produced close to the horizon is severely redshifted and most of it is captured rather than escaping to a distant observer. As a result, the radiative output is limited, and the efficiency cannot exceed the standard value determined by the specific energy at the innermost stable circular orbit, $\eta\simeq 5.7\%$.
In contrast, for a naked singularity, the absence of an event horizon allows the released binding energy to escape to infinity. In such a case, the theoretical radiative efficiency can, in principle, approach $100\%$ if the infalling matter loses its entire rest–mass energy before reaching the singularity. Here we consider the Joshi-Malafarina-Narayan(JMN) type 1 NaS psJoshi1 .
The line element of a general spherically symmetric, static spacetime is given as follows (we discuss it in detail in Appendix B):
@@
In the strong-field regime of general relativity, tidal forces provide a physically meaningful and observer-independent probe of spacetime curvature. These effects are most naturally described using the covariant geodesic deviation equation, which encodes the relative acceleration of nearby worldlines directly in terms of the Riemann tensor. The behavior of relativistic tidal stresses has been extensively investigated in black hole and compact-object spacetimes, revealing their sensitivity to both local curvature and global spacetime structure Marck:1983 ; Ishii:2005 ; Chicone:2005 . As such, tidal observables offer a robust framework for diagnosing extreme gravitational environments beyond purely causal or kinematical considerations. We investigate the effect of tidal forces in a static spherically symmetric (SSS) spacetime. To explore tidal forces in the framework of general relativity, we analyze the geodesic deviation equation:
@@
where $R^{\mu}_{\nu\rho\sigma}$ and $v^{\nu}$ denote the Riemann curvature tensor and the unit tangent vector of the geodesic, respectively, and $\eta^{\mu}$ represents the geodesic deviation vector. In the presence of strong energy conditions and non-zero curvature, each point on the test body follows a unique geodesic, leading to stretching and squeezing, known as tidal effects. To quantify these effects, which measures the distance between infinitesimally close geodesics. The tetrad basis related to the freely falling frame is expressed as:
@@
where $\eta_{\hat{\mu}\hat{\nu}}=\text{diag}(-1,1,1,1)$ represents components of the Minkowski metric. The hat indices denote the tetrad basis, while those without denote the coordinate basis. Additionally, $\hat{e}^{\mu}_{0}=v^{\mu}$, implying that the tetrad basis equals the 4-velocity vector of the observer when timelike zhang2018tidal . The unit vectors $\{\hat{e}^{\mu}_{1},\hat{e}^{\mu}_{2},\hat{e}^{\mu}_{3}\}$ correspond to orthogonal spatial directions in the observer’s frame chandrasekhar1983mathematical . Using the tetrad basis from Eq. (7), the geodesic deviation vector (or separation vector) can be expanded as:
@@
The spread in specific orbital energies imparted to the stellar material during a tidal encounter primarily reflects the differential gravitational potential across the star evaluated at the tidal radius, under the assumption that the star retains approximate hydrostatic balance prior to disruption (see Lacy et al. 1982 Lacy:1982 ). In this picture, fluid elements on the near and far sides of the star gain and lose orbital energy respectively due to the tidal field, producing the characteristic energy distribution that governs fallback rates after disruption.
A significantly broader dispersion of internal energies characterizes the star following the encounter, with some fluid having a positive energy (and thus staying unbound) and some having a negative energy (and hence being bound to the BH). The fluid elements move in Keplerian orbits. Following a Keplerian period T, the bound debris return near the pericenter and are connected to their (negative) energy E by:
@@
The fallback rate implied by Eq. (16) is often associated with the canonical $\dot{M}\propto t^{-5/3}$ scaling. It is important to note, however, that this behaviour is strictly asymptotic. While the early-time fallback can deviate significantly from a pure power law—particularly during the return of the most bound debris—the late-time behaviour approaches the $t^{-5/3}$ scaling provided that a nonzero amount of material occupies orbits near $E=0$, as is evident from Eq. (16). This asymptotic behaviour does not require a perfectly flat $dM/dE$, but follows generically once debris with small specific energies is present.
From an astrophysical perspective, tidal effects play a central role in a variety of high-energy phenomena, most notably tidal disruption events (TDEs), where stars experience extreme relativistic tidal stresses during close encounters with compact objects. The fallback dynamics, luminosity evolution, and observational signatures of TDEs are known to be highly sensitive to the strong-field spacetime geometry and relativistic corrections to stellar motion Guillochon:2013 ; Tejeda:2017 . This sensitivity makes tidal observables a promising bridge between theoretical models of strong-field gravity and astrophysical data, particularly in regimes where horizonless compact objects or nonstandard spacetime geometries may be relevant. Deviations from the canonical fallback rate can arise in cases of partial disruption or when the stellar internal structure plays an important role, as demonstrated in both analytical and numerical studies Lacy:1982 ; Lodato:2009 . These effects primarily influence the early-time evolution of the fallback rate, while the late-time behaviour remains consistent with the $t^{-5/3}$ scaling. Observational studies have also tested this prediction by comparing measured tidal disruption event light curves with theoretical fallback models, finding broad consistency with the asymptotic $t^{-5/3}$ behaviour fallowing for significant early-time deviations.
The efficient r-process nucleosynthesis observed in GW170817 highlights the crucial role of tidal disruption and mass ejection in heavy-element production. The suppression of comparable r-process signatures in neutron star–black hole mergers indicates that horizon absorption limits ejecta formation, whereas horizonless compact objects may sustain strong tidal interactions capable of enhancing mass ejection and nucleosynthesis, even at high total masses.
@@ II Tidal disruption radius in BHs and naked singularities
The way stars are destroyed close to supermassive compact objects, which might not be typical black holes but instead have scalar ”hair” or even be naked singularities is examined in Andre:2024bia . The authors demonstrate how scalar fields can significantly increase tidal forces in the vicinity of black holes. At the same time, flares may be powered by stellar collisions with a dense ”gray shell” in the case of a naked singularity. According to their findings, tidal disruption events might provide a means of differentiating between black holes and other, more unusual central objects in galaxies. The tidal disruption radius is calculated on the principle of the radial tidal force exerted on the test particle, Eq. (12)(here we consider a NS with some solar mass), and the gravity on the surface of the test particle is equal. However, assuming that the tidal disruption of the star takes place when the self-gravitational attraction on its surface equals the tidal force, which is an elementary but useful approximation for the tidal disruption radius,
$\mid{T_{R}}\mid R_{n}=\frac{Gm}{R_{n}^{2}}.$ (17)
Equation (17) is adopted from previous studies of tidal disruption in black hole and naked singularity spacetimes Andre:2024bia and should be understood as a heuristic criterion indicating the onset of significant tidal deformation, rather than as a fully relativistic tidal radius. In a proper relativistic treatment, tidal effects are characterized by the eigenvalues of the tidal tensor constructed from the Riemann curvature projected onto the local orthonormal frame of the neutron star, and quantities such as the stellar radius cannot be interpreted as coordinate distances. Fully relativistic analyses have shown that the tidal disruption threshold depends sensitively on the neutron star compactness, equation of state, and the spacetime curvature of the central object Thorne:1987 ; Hinderer:2008 ; Beloborodov:1992 ; Kesden:2012 ; Stone:2013 . Our use of Eq. (17) is therefore intended to provide qualitative insight into the relative strength of tidal effects in naked singularity spacetimes compared to black holes, rather than a precise or invariant disruption criterion. Here, the radial tidal force component is,
@@ II.2 Roche limit in JMN1 NaS spacetime
Fig. (1) demonstrates tidal disruption radius in JMN1 NaS for $M_{0}<2/3$. As shown in Eq. (24), a real positive value of $T_{R}$ is when $M_{0}<2/3$, which implies that tidal disruption is because of the stretching forces while $M_{0}>2/3$ compressive forces are present. Tidal deformation is also closely related to the Jacobi fields that are fully explained in Appendix C.
@@
The evolution of flares with time near a supermassive compact object or the merger of two compact binaries arises due to the disruption of the compact object. Here, we analytically show the difference between naked singularities and BHs, relating the light curve to the internal density distribution of the star. As it has been previously shown Lodato:2009 , the standard light curve is proportional to $t^{-5/3}$ only holds at late times. Unbound debris always carries larger mechanical energies than the rest mass energy.
@@
As it is shown in the strong gravity correction order $1/c^{2}$, $dE/d\tau>0$. From Fig.(1), we can conclude that tidal disruption within the ISCO to the event horizon is possible within the stellar mass mergers of binary BH and NS ($1.5M_{\odot}$ to $7.3M_{\odot}$). In the Schwarzschild case, the innermost unstable circular orbit (IUCO) is $4M$. Hence, in between the $4M$ to $6M$ unstable circular orbit can be possible. If the tidal disruption radius is within this range, the tidal disruption radius is the periapsis of the NS. If a NS approaches a BH so closely that it is affected by tidal distortions and even disruption, it cannot be considered a “point mass” interacting with the BH. Such effects become significant when the ‘tidal radius’ is as small as the pericenter, $R_{T}<r_{min}$. Therefore, complete disruption requires $\beta=R_{T}/r_{min}<1$.
For JMN-1 NaS proper time for a circular orbit is given as follows:
$\tau=\frac{2\pi r\sqrt{R_{b}}}{\sqrt{GM}}\sqrt{1-\frac{3GM}{c^{2}R_{b}}}$ (29)